# Feature Hedging: Correlated Features Break Narrow Sparse Autoencoders

### v1

@@ A.2 Training details for LLM SAEs
A.3 Choice of hedging hyperparameter N
@@
A.3.1 Extending LLM SAEs
We train two versions of extension SAEs - one for L1 loss SAEs and one for BatchTopK SAEs. In both cases, we begin with a pretrained SAE and add $N$ latents randomly initialized with norm 0.1, and with the same encoder and decoder directions, following Olah et al. [18]. For the BatchTopK SAEs, we simply train the SAE from this point as normal, as the TopK auxiliary loss [11] will naturally ensure that the newly added latents do not simply die off.
@@ A.3.1 Extending LLM SAEs
A.4 Additional case study dashboards
@@
A.5 Toy balance matryoshka SAEs
@@
A.6 Toy unbalanceable matryoshka SAEs
@@
A.7 SAE evaluation
A.7.1 SAEBench evals
We evaluate our SAEs mainly using SAEBench [14]. All evals are performed using default settings. We run all evaluations on an Nvidia H100 GPU with 80gb GPU memory. We evaluate on the following SAEBench tasks:
@@ A.7.1 SAEBench evals
k-sparse probing builds on the work of Gurnee et al. [12], where the goal is to create a linear probe from model activations using only $k$ neurons as input to the probe. This was adapted for use as an SAE evaluation technique by Gao et al. [11]. We focus mainly on $k=1$ sparse probing, which means finding the single best SAE latent that serves as a classifier for a given concept, and evaluating the accuracy of that latent. SAEBench includes supervised classification datasets on which k-sparse probing is evaluated.
@@
The feature absorption metric in SAEBench is a variation on the metric defined in the original feature absorption work [5]. This metric uses a first-letter spelling task and first identifies the “main” latents for that task using k-sparse probing [12]. Then, the metric identifies cases where a linear probe is able to correctly perform the first-letter classification task, but the “main” SAE latents fail to perform the task. The metric also looks for other latents that project onto the linear probe direction and fire in place of the “main’ latents. Lower absorption is better.
@@
SCR builds on the SHIFT method from Marks et al. [17] to detect how well an SAE isolates concepts. The metric uses datasets where two properties are perfectly entangled, for instance “profession” and “gender”, and trains a biased probe on these concepts. The SCR metric then detects how well $k$ SAE latents can be ablated to de-bias the probe. If the SAE latents learn disentangled concepts, then it should only take a few SAE latents to perfectly de-bias the probe. A high SCR score means the SAE latents represent disentangled concepts.
@@
A.7.2 Parts of speech (POS) probing dataset
We are interested as well in general, high-frequency concepts that we expect should be learned in the inner-most levels of a matryoshka SAE. These concepts should be the most affected by both absorption and hedging, as these concepts can be considered parent concepts to most other concepts. Parts of speech (POS) is a great test-case for these general concepts, and are not covered by the SAEBench sparse probing task. As such, we create our own custom POS dataset using the Penn Treebank tagged sentences [16].
@@ A.7.2 Parts of speech (POS) probing dataset
We pass these tagged sentences through an LLM, and then collect activations for the final token of position of each word at a given layer in the LLM. We create a binary classification dataset for each of these parts of speech, and perform k-sparse probing [12] on SAE latents to find the top k latents that act as a classifier for each of these parts of speech.
A.7.3 Balance matryoshka SAE full results

### v2

@@ A.3 Training details for LLM SAEs
A.4 Choice of hedging hyperparameter N
@@
A.4.1 Extending LLM SAEs
We train two versions of extension SAEs - one for L1 loss SAEs and one for BatchTopK SAEs. In both cases, we begin with a pretrained SAE and add $N$ latents randomly initialized with norm 0.1, and with the same encoder and decoder directions, following Olah et al. (2024). For the BatchTopK SAEs, we simply train the SAE from this point as normal, as the TopK auxiliary loss (Gao et al., 2024) will naturally ensure that the newly added latents do not simply die off.
@@ A.4.1 Extending LLM SAEs
A.5 Case study: adding a new latent to an existing SAE
We next explore how hedging affects a real SAE. We trained a L1 SAE on Gemma-2-2b layer 12 with width 8192 for 250M tokens on the Pile (Gao et al., 2020), then add a new latent to the SAE, and continue training both the original SAE and the extended SAE for another 250M tokens.
0.3/css/bootstrap.min.
/bootstrap.min.css_integrity="sha3"
/bootstrap.min.css">____link
>____<link_rel="stylesheet"_type="
8"/>< link_rel=stylesheet_href="..//doc
png"><link_rel="manifest"_href="
(b) Latent 3094, which had the largest negative $\delta$-projection after adding latent 8192. This latent tracks ‘‘rel’’ in HTML, used for CSS in HTML.
Figure 14: Sample top activating examples for case study latents.
A.6 Additional case study dashboards
@@
A.7 Toy balance matryoshka SAEs
@@
A.8 Toy unbalanceable matryoshka SAEs
@@
A.9 SAE evaluation
A.9.1 SAEBench evals
We evaluate our SAEs mainly using SAEBench (Karvonen et al., 2025). All evals are performed using default settings. We run all evaluations on an Nvidia H100 GPU with 80gb GPU memory. We evaluate on the following SAEBench tasks:
@@ A.9.1 SAEBench evals
k-sparse probing builds on the work of Gurnee et al. (2023), where the goal is to create a linear probe from model activations using only $k$ neurons as input to the probe. This was adapted for use as an SAE evaluation technique by Gao et al. (2024). We focus mainly on $k=1$ sparse probing, which means finding the single best SAE latent that serves as a classifier for a given concept, and evaluating the accuracy of that latent. SAEBench includes supervised classification datasets on which k-sparse probing is evaluated.
@@
The feature absorption metric in SAEBench is a variation on the metric defined in the original feature absorption work (Chanin et al., 2024). This metric uses a first-letter spelling task and first identifies the “main” latents for that task using k-sparse probing (Gurnee et al., 2023). Then, the metric identifies cases where a linear probe is able to correctly perform the first-letter classification task, but the “main” SAE latents fail to perform the task. The metric also looks for other latents that project onto the linear probe direction and fire in place of the “main’ latents. Lower absorption is better.
@@
SCR builds on the SHIFT method from Marks et al. (2025) to detect how well an SAE isolates concepts. The metric uses datasets where two properties are perfectly entangled, for instance “profession” and “gender”, and trains a biased probe on these concepts. The SCR metric then detects how well $k$ SAE latents can be ablated to de-bias the probe. If the SAE latents learn disentangled concepts, then it should only take a few SAE latents to perfectly de-bias the probe. A high SCR score means the SAE latents represent disentangled concepts.
@@
A.9.2 Parts of speech (POS) probing dataset
We are interested as well in general, high-frequency concepts that we expect should be learned in the inner-most levels of a matryoshka SAE. These concepts should be the most affected by both absorption and hedging, as these concepts can be considered parent concepts to most other concepts. Parts of speech (POS) is a great test-case for these general concepts, and are not covered by the SAEBench sparse probing task. As such, we create our own custom POS dataset using the Penn Treebank tagged sentences (Marcus et al., 1993).
@@ A.9.2 Parts of speech (POS) probing dataset
We pass these tagged sentences through an LLM, and then collect activations for the final token of position of each word at a given layer in the LLM. We create a binary classification dataset for each of these parts of speech, and perform k-sparse probing (Gurnee et al., 2023) on SAE latents to find the top k latents that act as a classifier for each of these parts of speech.
A.9.3 Balance matryoshka SAE full results